OpenStudy (anonymous):

How many solutions will this system of equations have? y = 3.5x - 3.5 y = -3.5x + 3.5 No solution Infinite solutions One solution Two solutions

OpenStudy (cp9454):

infinite solutions

OpenStudy (imstuck):

Why do you say infinite?

OpenStudy (imstuck):

Infinite is wrong...just sayin.

OpenStudy (cp9454):

When you don't know the rules just don't say it. It is correct.

OpenStudy (imstuck):

why do you say infinite? When you solve for x you get 1.

OpenStudy (imstuck):

3.5x-3.5=-3.5x+3.5 3.5x+3.5x=3.5+3.5 7x=7 x=1

OpenStudy (anonymous):

So, how would it be no solution?

OpenStudy (cp9454):

look it's like\[\frac{ a1 }{ a2 } = \frac{ b1 }{ b2 } = \frac{ c1 }{ c2 }\] then the pair of equations have infinite many solutions

OpenStudy (imstuck):

so now we have no solution, infinite solutions, but you can solve it for x. I don't understand how you can solve it for x and there is not a solution. you need to explain that please.

OpenStudy (anonymous):

So, what is the REAL answer to this problem?

OpenStudy (cp9454):

oh, yes it,s unique solution. i m really sorry. as here \[\frac{ a1 }{ a2 } \neq \frac{ b1 }{ b2 }\] i didn't see it really sorry.

OpenStudy (imstuck):

So it does have 1 solution, then?

OpenStudy (imstuck):

x=1...one solution. I'm sticking with that. Don't see how I am incorrect.

OpenStudy (anonymous):

Alright I'll let you two know.

OpenStudy (anonymous):

One solution is correct!

OpenStudy (imstuck):

Thank you...for the confirmation and the medal.

OpenStudy (imstuck):

I'm not a math teacher for nothing....